OpenStudy (anonymous):

Simplify 12 to the 3rd over 12 to the 7th. 12^4 1 over 12 to the 4th 1 over 12 to the negative 4th 12^10

OpenStudy (anonymous):

@phi

OpenStudy (phi):

\[ \frac{12^3}{12^7} \] any idea what \[ \frac{12}{12} \] is ?

OpenStudy (anonymous):

12/12 would equal to 1 right?

OpenStudy (phi):

yes. as you know , when you multiply fractions you multiply top times top and bottom times bottom, right ?

OpenStudy (anonymous):

ya top and top bottom and bottom

OpenStudy (phi):

\[ \frac{12\cdot 12\cdot 12}{12\cdot 12\cdot 12} \cdot \frac{1}{12\cdot 12\cdot 12\cdot 12}\] do you see that is 12^3 / 12^7 ?

OpenStudy (anonymous):

yes

OpenStudy (phi):

do you also see the first part can be simplified? something divided by itself is what ?

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

i dont see what can be divided

OpenStudy (phi):

or perhaps you should notice \[ \frac{12\cdot 12\cdot 12}{12\cdot 12\cdot 12} = \frac{12}{12} \cdot \frac{12}{12} \cdot \frac{12}{12}\]

OpenStudy (anonymous):

mmmm

OpenStudy (anonymous):

if they got multiplied together and divided? maybe??

OpenStudy (phi):

the right side, if you multiply tops you get 12*12*12 and ditto for the bottom. both sides mean the same thing but what is 12/12 if you simplify first?

OpenStudy (anonymous):

12/12 simplified is 1

OpenStudy (phi):

and what is 1*1*1? it is 1 \[ \frac{12\cdot 12\cdot 12}{12\cdot 12\cdot 12} = \frac{12}{12} \cdot \frac{12}{12}\cdot \frac{12}{12} = 1 \cdot 1 \cdot 1 \] (12*12*12)/(12*12*12) is an example of something (in this case 12*12*12) divided by itself. you get 1

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

ok

OpenStudy (phi):

so simplify \[ \frac{12\cdot 12\cdot 12}{12\cdot 12\cdot 12} \cdot \frac{1}{12\cdot 12\cdot 12\cdot 12} \]

OpenStudy (anonymous):

1/12^4?

OpenStudy (anonymous):

because the 4 12?

OpenStudy (phi):

and the first part (12*12*12)/(12*12*12) is just 1 and when we multiply by 1 we can ignore it. so yes 1/12^4

OpenStudy (anonymous):

alrighty:D

OpenStudy (anonymous):

and im taking it as 1/12^4 is the answer also?

OpenStudy (phi):

the short cut is subtract top exponent minus bottom exponent \[ \frac{12^3}{12^7} = 12^{3-7} = 12^{-4} = \frac{1}{12^4} \]

OpenStudy (anonymous):

ohhhh

OpenStudy (anonymous):

let me take a pic of thats because thats useful

OpenStudy (anonymous):

ok thats it

OpenStudy (anonymous):

let me submmit it

OpenStudy (anonymous):

your the best we got a 100%

OpenStudy (anonymous):

thanks man your the best if i could rate you id give u a million stars xD

OpenStudy (phi):

yw

OpenStudy (anonymous):

OK THANKS MAN BYE!!